# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

### v2

@@ 5 Conclusion
The proposed AGI test bed provides a systematic approach to AGI evaluation through 12 essential tests. By starting from an unbiased state and focusing on fundamental input-driven learning, this test bed aligns with contemporary neuroscience insights.
AGITB can be solved by humans but remains unsolvable by classical algorithms and current state-of-the-art AI. This gap between human and machine performance serves as strong empirical evidence that AGITB is a meaningful benchmark for evaluating AGI capabilities. While no known computational system—including deterministic algorithms and advanced neural networks—has successfully solved it, this alone does not constitute formal proof of its fundamental difficulty. However, the fact that humans can solve AGITB while all existing AI systems fail, suggests that it captures a crucial aspect of general intelligence.

### v3

@@ 4 Conclusion
Unlike conventional benchmarks that evaluate high-level task performance, such as question answering or language translation, AGITB assesses whether a system exhibits behaviour consistent with the core operational principles of the biological cortex. It focuses on low-level computational properties that are both biologically grounded and essential for the emergence of general intelligence. The proposed testbed introduces a systematic framework for evaluating AGI through twelve fundamental tests. By requiring models to begin from an unbiased initial state and adapt solely through exposure to structured or unstructured input, AGITB reflects key insights from contemporary neuroscience regarding input-driven learning and cortical plasticity.
@@
AGITB remains solvable by humans but unsolved by classical algorithms and current state-of-the-art AI systems. This persistent performance gap provides strong empirical evidence that AGITB captures essential properties of general intelligence. While the absence of a computational solution does not formally prove its adequacy as a benchmark, the consistent success of humans where machines fail suggests that AGITB not only distinguishes narrow from general intelligence, but also offers a valuable framework for guiding the development of truly general AI systems.